OpenStudy (anonymous):

Prove the following trigonometric identity using the double angle formula. [sin(2x)/sinx] - [cos(2x)/cosx] = secx

OpenStudy (freckles):

Try to write left hand side as term do that by first combining the fractions

OpenStudy (freckles):

as one term*

OpenStudy (anonymous):

Make it into [sin(2x)cosx - cos(2x)sinx/sinxcosx]...

OpenStudy (anonymous):

Right...

OpenStudy (freckles):

right use sin difference identity on top

OpenStudy (freckles):

\[\sin(a-b)=\sin(a)\cos(b)-\sin(b)\cos(a)\]

OpenStudy (anonymous):

? I'm not getting where the difference identity comes into play.

OpenStudy (anonymous):

There isn't any sin(a+b) or sin(a-b)

OpenStudy (freckles):

\[\frac{\sin(2x)\cos(x)-\sin(x)\cos(2x)}{\sin(x)\cos(x)}\]

OpenStudy (freckles):

I see sin(a)cos(b)-sin(b)cos(a) on top...

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Holy pellet! Yeah! Didn't catch that...

OpenStudy (anonymous):

So sin (2x - x)/sinxcosx...

OpenStudy (freckles):

2x-x=x so we have sin(x)/(sin(x)cos(x))

OpenStudy (freckles):

something cancels

OpenStudy (freckles):

And I hope you see that we have sec(x)

OpenStudy (anonymous):

Wow yeah. Then just the sec defintion

OpenStudy (freckles):

This wasn't the only approach you could have taken

OpenStudy (anonymous):

*definition

OpenStudy (freckles):

\[\frac{\sin(2x)}{\sin(x)}-\frac{\cos(2x)}{\cos(x)} \\ \frac{2 \sin(x) \cos(x)}{\sin(x)}-\frac{\cos^2(x)-\sin^2(x)}{\cos(x)} \\ 2\cos(x)-(\cos(x)-\frac{\sin^2(x)}{\cos(x)} ) \\ 2\cos(x)-\cos(x)+\frac{\sin^2(x)}{\cos(x)} \\ \cos(x)+\frac{\sin^2(x)}{\cos(x) } \\ \frac{\cos^2(x)+\sin^2(x)}{\cos(x)} \\ \frac{1}{\cos(x)} \\ \sec(x)\]

OpenStudy (freckles):

But I like the first approach better

OpenStudy (freckles):

1st) Use double angle identities on both sin(2x) and cos(2x) 2nd) Canceled the sin's in the first fraction and wrote the second fraction as two fractions in my mind while canceling a cos in the first fraction of the rewriting of the second fractions 3rd) Distribute the negative in front of the second part there 4th) Combined like-terms 5th) wrote it one term by forcing a common denominator 6th) Used a Pythagorean identity

OpenStudy (freckles):

First way much prettier

OpenStudy (anonymous):

My thanks